# TuRTLe: A Unified Evaluation of LLMs for RTL Generation

### v1

@@ Abstract
The rapid advancements in LLMs have driven the adoption of generative AI in various domains, including Electronic Design Automation (EDA). Unlike traditional software development, EDA presents unique challenges, as generated RTL code must not only be syntactically correct and functionally accurate but also synthesizable by hardware generators while meeting performance, power, and area constraints. These additional requirements introduce complexities that existing code-generation benchmarks often fail to capture, limiting their effectiveness in evaluating LLMs for RTL generation. To address this gap, we propose TuRTLe, a unified evaluation framework designed to systematically assess LLMs across key RTL generation tasks. TuRTLe integrates multiple existing benchmarks and automates the evaluation process, enabling a comprehensive assessment of LLM performance in syntax correctness, functional correctness, synthesis, PPA optimization, and exact line completion. Using this framework, we benchmark a diverse set of open LLMs and analyze their strengths and weaknesses in EDA-specific tasks. Our results show that reasoning-based models, such as DeepSeek R1, consistently outperform others across multiple evaluation criteria, but at the cost of increased computational overhead and inference latency. Additionally, base models are better suited in module completion tasks, while instruct-tuned models perform better in specification-to-RTL tasks.
@@ I Introduction
The rapid advancements in large language models (LLMs), known for their ability to process and generate human-like text, have unlocked new possibilities across a wide range of domains [1, 2]. Domain-specific LLMs have gained significant attention due to their strong performance in specialized tasks, including financial engineering [3], biomedical research [4], and scientific computing [5, 6]. LLMs also assist developers by suggesting code snippets, solving common coding challenges, and providing clear explanations of complex concepts [7, 8, 9].
In the field of Electronic Design Automation (EDA), researchers are increasingly exploring the use of LLMs to accelerate hardware design[10, 11, 12]. In the traditional digital system design flow, engineers invest significant effort in implementing precise functionality using hardware description languages (HDLs). LLM-based solutions aim to bridge this gap by translating functional specifications directly into HDL code, such as Verilog. This approach has the potential to revolutionize hardware design and verification by streamlining Verilog coding, optimizing circuit implementations, and automating time-consuming design tasks [13]. Recent research has explored various techniques for LLM-driven Verilog code generation, including: prompt engineering to enhance model responses [14], training and fine-tuning on Verilog-specific datasets [15, 16, 17, 18, 19], and agent-based approaches to improve iterative refinement and debugging [20, 21, 22, 23, 24]. However, several research questions about the quality and fidelity of the generated Verilog code and, hence, the hardware design, remain largely unanswered. As for code generation for other programming languages, LLM-based solutions need to generate code that is syntactically and functionally correct, i.e., , the generated code must compile and generate the correct results. High-quality code must also provide reasonable performance. EDA tools pose the additional constraints that hardware generators need to be able to take the generated RTL and be able to synthesize it to produce correct hardware components. Additionally, non-functional constraints, such as area and power, play an important role when resources are limited. It is evident that evaluation methodologies and benchmarks designed to assess LLM that produce traditional code do not fully cover the entire evaluation space and are not completely suitable to evaluate LLM-based EDA.
@@
This work introduces TuRTLe, a comprehensive evaluation framework for RTL generation. It integrates multiple benchmarks into a single fully-automated evaluation framework, providing a standardized and comprehensive assessment of code. Coverage is provided for different RTL design goals, including syntax and correctness of the generated code, synthesizability of the hardware circuits produced by hardware generators, together with their corresponding performance, area, and power metrics. These goals are tested under different conditions, in particular single-line completion, module completion, and generation from specifications expressed in natural language. TuRTLe is used to benchmark 13 models, providing a comprehensive, wide and updated review on the state of open models. This includes general purpose models, coding, and RTL-specific LLMs.
Our analysis show that temperature, context length, model structure (base vs. instruct), and reasoning mechanisms all play a crucial role in LLM performance across EDA-related tasks. Base models demonstrate stronger performance in module completion, effectively filling in partial RTL definitions, while instruct-tuned models excel at generating RTL from natural language specifications, leveraging their training in instruction-following tasks. Models employing reasoning prompting chains, such as DeepSeek R1, exhibit remarkable efficiency in handling complex tasks like specification-to-RTL, though at the cost of increased inference time and token generation. Overall, our findings indicate that while current LLMs are highly proficient in generating syntactically correct RTL code, they struggle with functional correctness. However, once functional correctness is achieved, synthesizability is generally not an issue, and the resulting circuit designs are often comparable to human-defined modules.
@@ II Evaluation Methodology
Evaluating LLM-generated RTL code requires a comprehensive and systematic approach that considers the many aspects of the hardware design process. TuRTLe integrates existing benchmarks including VeriGen [15], VerilogEval [25, 26], RTL-Repo [27], and RTLLM [28], within a single infrastructure, allowing for direct comparison of LLM performance across various RTL generation tasks. Focusing on versatility and automation in the construction and running of TuRTLe, we choose a widely adopted framework in the field of code generation (BigCode Evaluation Harness [29]) as a foundation. This tool can easily load a variety of generative models, and already implements multiple LLM tasks and metrics. Focusing on scalability, we fork from the vLLM Code Harness project [^1] https://github.com/iNeil77/vllm-code-harness, a particularly efficient derivative, which facilitates running larger models through the vLLM [30] library. Through these choices, TuRTLe is compatible by design with most popular practices in the field.
This section first reviews three tasks involved in RTL generation that TuRTLe focuses on: single-line completion, module completion, and specification-to-RTL conversion, all in §II-A. Next, it outlines the five RTL design goals assessed by TuRTLe: line-level contextual accuracy, syntax correctness, functional correctness, synthesizability, and post-synthesis quality (see §II-B). To evaluate LLM performance on these goals, we employ a set of metrics, including exact matching, Pass@1, and a novel PPA-Score (detailed in §II-C). Table I shows the relationships between the various tasks, design goals, metrics, and tools used in the TuRTLe framework.
@@ II-A Generation Tasks
Our evaluation framework assesses LLMs across three fundamental RTL code generation tasks, each representing a distinct levels of complexity and scale. The smallest of tasks, in the sense of generation length, is Single-Line Completion (SLC), which focuses on the model’s ability to predict the next line of code given a partial context. This task closely resembles auto-completion scenarios, where engineers rely on coding assistants to streamline development, and evaluates the capacity of models to produce contextually coherent code.
@@
Finally, the third and most complex task is Specification-to-RTL (S2R) generation, where the LLM must generate a complete implementation from a natural language hardware specification. This task does not provide a predefined module interface, so models must infer additional information from the specification, such as the module name, the correct ports and parameter types and names. This mimics the process of translating human-readable specifications into working hardware descriptions. Given their fundamental differences in complexity and nature, results for these three approaches are reported separately.
@@ II-B Design Goals
TuRTLe defines five key design goals to evaluate LLM-generated RTL code, ensuring a comprehensive assessment of validity, correctness, feasibility, and efficiency within the hardware design workflow. These goals are aligned with the supported tasks to maintain compatibility. Given SLC, MC and S2R, we consider the following design goals.
Line-level Contextual Accuracy (LCA) tests the ability of the models to produce line completions that are as close as possible to a given reference answer. This goal aligns with the SLC task, and evaluates whether the predicted line is both syntactically correct and contextually coherent within the surrounding logic.
The four remaining goals can all be applied to both MC and S2R tasks. Syntax Correctness (STX) ensures that the generated HDL code complies with the language’s grammar rules and can be processed by standard HDL analyzers, synthesizers, and simulators [31]. The correctness of generated designs is verified by parsing and elaboration, using tools that check for syntax errors, missing constructs, and structural inconsistencies. Functional Correctness (FNC) ensures that the generated HDL code exhibits the expected behavior as specified in the design requirements (prompt) [32, 33]. FNC is verified through behavioral simulation, where the generated design is executed alongside a predefined testbench in an HDL simulator. While this approach provides a practical verification method, it relies on the assumption that testbenches accurately capture functional discrepancies.
Synthesizability (SYN) assesses whether the generated HDL code can be successfully synthesized into a gate-level netlist using a synthesis tool. In practical hardware design, synthesis tools support only a subset of HDL constructs, making synthesizability a key criterion for real-world applicability. SYN is verified by analyzing and synthesizing the generated RTL code using a synthesis toolchain, ensuring that the design is implementable in hardware. This is critical for models that generate RTL intended for manufacturable chip designs.
Finally, Post-Synthesis Quality (PSQ) evaluates the implementation efficiency of the synthesized design based on Power, Performance, and Area (PPA). PPA metrics are widely used in chip design optimization to compare different implementations and ensure that a design meets key hardware constraints [34]. In this work, we use PPA to compare LLM-generated designs with manually optimized reference implementations. We extract area and power directly from the PPA report and evaluate performance based on maximum delay, defined as the difference between the clock period and the worst slack reported by static timing analysis. In this way, all three PPA metrics are represented as positive numbers, where 00 represents the minimum possible value, leaving the maximum unbounded.
TABLE I: Design goals and metrics for benchmarking RTL generation. The use of open-source tools ensures broader accessibility and reproducibility.
@@
SYN OpenLANE (Yosis)
PSQ PPA-Score OpenLANE (OpenROAD)
@@ II-C Measures and Metrics
To evaluate the five goals described above (LCA, STX, FNC, SYN and PSQ) we use different tools and evaluation measures tailored to each case. For LCA, prior work [27] has explored both exact and fuzzy matching techniques. Exact Matching (EM) determines whether the generated line precisely matches the reference HDL line in the dataset, while Fuzzy Matching relies on edit distance and semantic similarity metrics. Our experimentation shows both metrics are strongly correlated, which is why we report the more precise EM alone.
For MC and S2R tasks, we set up an evaluation pipeline that sequentially tests STX, FNC, SYN, and PSQ. STX is evaluated by compiling a design along with its testbench using Icarus Verilog [35] and checking for errors. If no errors are issued at compile time, FNC is evaluated by running the simulation executable generated by the compiler and checking if the testbench passes. Functionally correct codes are tested for SYN, by elaborating the design with OpenLANE (Yosys) [36]. For the four previous goals (LCA, STX, FNC, and SYN), we report the same metric Pass@$1$. This is a Pass@$k$ metric [37, 38], which measures the probability that at least one of $k$ generated solutions passes the corresponding test criteria. We set $k=1$ to focus on the requirement of generating the correct result on the first try.
For those designs that pass SYN, PSQ is computed by using OpenLANE [39] to synthesize the code into a netlist and extract post-synthesis PPA metrics. Designs are synthesized using the SKY130A open-source PDK [40] with a $10ns$ delay constraint. In this work, we introduce a novel metric called PPA-score to measure PSQ, focusing on the comparison between LLM-generated synthesizable code and a golden, human-crafted reference.
@@
Given a benchmark with $n$ problems, for each problem $i\in\{1,\dots,n\}$, we generate $m$ candidate solutions. Each generation $j\in\{1,\dots,m\}$ is processed through the evaluation pipeline sequentially: STX, FNC, SYN, PSQ. Each stage only processes results that passed the previous one, creating a cascade score where $STX\geq FNC\geq SYN\geq PSQ$ (failures in previous stages are reported as automatic fails in the next ones). STX, FNC and SYN are binary evaluations, pass or fail, which are aggregated using Pass@$k$. However, PSQ requires not only to aggregate numeric values that have to be analyzed in comparison with the reference PPA of the golden solution, but also to take into account that models that produce more synthetizable code can be evaluated on a larger set of problems, thus increasing the challenge and confidence in the results obtained. To achieve this, let $p_{i,j}$ represent the PPA metric (power, performance, or area) from the LLM for candidate $j$ of problem $i$:
@@
    Each $p_{i,j}$ is compared against the corresponding PPA value $g_{i}$ of the golden solution. For that, instead of aggregating $p_{i,j}$ we compute $\hat{p}_{i,j}=p_{i,j}/g_{i}\in(0,+\infty)$.
@@
    For generations that do not pass STX, FNC and SYN evaluations, $p_{i,j}$ cannot be computed. As a result, we set a failure value of $\hat{p}_{i,j}=2$ (e.g. producing a design two times bigger than the human reference in the case of the area metric). This approach also requires us to limit results that pass the previous evaluations but perform worse than this threshold (i.e. $\hat{p}_{i,j}>2$).
@@
Following these steps, if a design has score = 0 means it requires twice the area, the power or the performance, when compared to the human reference. Any measurement worse than double (i.e., on the negative side of the score) is clipped to zero (because we set $\hat{p}_{i,j}=2$). A score = 0.5 are designs with an area, power or performance equal to that of the human reference. Finally, a score = 1 can only be obtained by chips which occupy no space, execute in no time, and consume no energy. The final formula considering all generations of an LLM for a given benchmark is then computed as:

### v2

@@ Abstract
Rapid advancements in LLMs have driven the adoption of generative AI in domains like Electronic Design Automation (EDA). Within the field of software development, EDA presents unique challenges derived from specific requirements of generated RTL code; RTL code must not only be syntactically correct and functionally accurate, but also synthesizable by hardware generators, while matching performance, power and area (PPA) constraints. These additional requirements introduce complexities that existing code-generation benchmarks often fail to capture, limiting their effectiveness in evaluating LLMs for RTL generation. To address this gap, we propose TuRTLe, a unified evaluation framework designed to systematically assess LLMs across key RTL generation tasks. TuRTLe integrates multiple existing benchmarks and automates the evaluation process, enabling a comprehensive assessment of LLM performance in syntax correctness, functional correctness, synthesis, PPA optimization, and exact line completion. Using this framework, a diverse set of forty open LLMs are assesed, tracking their strengths and weaknesses in EDA-specific tasks. Our results identify the best match for specific tasks (e.g., base models are better in module completion tasks, instruct-tuned models are better in specification-to-RTL tasks), while finding that recent models with autoregressive reasoning chain perform the best overall. Some benchmarks, particularly within syntax correctness, show signs of saturation, while others remain as open problems for LLMs.
@@ I Introduction
Advancements in large language models (LLMs) have unlocked new possibilities across a wide range of domains [1, 2]. Domain-specific LLMs have gained significant attention due to their strong performance in specialized tasks, including financial engineering [3], biomedical research [4], and scientific computing [5, 6]. In software related tasks, LLM can suggest code snippets, solve common coding challenges, and provide explanations of complex concepts [7, 8, 9, 10].
In the field of Electronic Design Automation (EDA), researchers are increasingly exploring the use of LLMs to accelerate hardware design[11, 12, 13]. In the traditional digital system design flow, engineers must implement precise functionality using hardware description languages (HDLs). LLM-based solutions aim to bridge this gap by translating functional specifications directly into HDL code, such as Verilog. This approach has the potential to revolutionize hardware design and verification by streamlining Verilog coding, optimizing circuit implementations, and automating time-consuming design tasks [14]. Recent research has explored various techniques for Verilog code generation, including prompt engineering to enhance model responses [15], training or fine-tuning on Verilog-specific datasets [16, 17, 18, 19, 20], and agent-based approaches to improve iterative refinement and debugging [21, 22, 23, 24, 25]. However, several research questions about the quality and fidelity of the generated Verilog code remain largely unanswered. LLM-based solutions need to generate code that is syntactically and functionally correct, but in EDA the generated code must also be synthesizable to produce correct hardware components, and achieve reasonable performance in non-functional indicators such as area and power.
@@
This work introduces TuRTLe, a comprehensive evaluation framework for RTL generation. It integrates multiple benchmarks into a single fully-automated evaluation framework, providing a standardized and comprehensive assessment of code. Coverage is provided for different RTL design goals, including syntax and correctness of the generated code, synthesizability of the hardware circuits produced by hardware generators, together with their corresponding performance, area, and power metrics, using a novel metric based on comparison against human performance. These goals are tested under different conditions, such as module completion, and generation from specifications expressed in natural language. TuRTLe is used to benchmark forty models, providing a comprehensive, wide and updated review on the state of open models. This includes general purpose, coding, and RTL-specific LLMs, as well as base, instruct and reasoning models, showing the strengths and weaknesses of each alternative.
Our analysis shows that temperature, context length and training strategy (e.g., base, instruct, reasoning) all play a crucial role in LLM performance across EDA-related tasks. Base models demonstrate stronger performance in module completion, effectively filling in partial RTL definitions, while instruct-tuned models are better at generating RTL from natural language specifications. Models employing reasoning prompting chains, exhibit remarkable performance across the board at the cost of increased inference time and token generation. Overall, our findings indicate that while current LLMs are highly reliable in generating syntactically correct RTL code, they still struggle with functional correctness and in achieving human-level non-functional indicators.
@@ II Evaluation Methodology
Evaluating LLM-generated RTL code requires a comprehensive and systematic approach that considers the many aspects of the hardware design process. TuRTLe integrates existing benchmarks including VeriGen [16], VerilogEval [26, 27], RTL-Repo [28], and RTLLM [29], within a single infrastructure, allowing for direct comparison of LLM performance across various RTL generation tasks. Focusing on versatility and automation in the construction and running of TuRTLe, we choose a widely adopted framework in the field of code generation (BigCode Evaluation Harness [30]) as a foundation. This tool can easily load a variety of generative models, and already implements multiple LLM tasks and metrics.
This section first reviews three tasks involved in RTL generation that TuRTLe focuses on: single-line completion, module completion, and specification-to-RTL conversion, all in §II-A. Next, it outlines the five RTL design goals assessed by TuRTLe: line-level contextual accuracy, syntax correctness, functional correctness, synthesizability, and post-synthesis quality (see §II-B). To evaluate LLM performance on these goals, we employ several metrics, including exact matching, Pass@1, and a novel PPA-Score (detailed in §II-C). Table I shows the relationships between the various tasks, design goals, metrics, and tools used in the TuRTLe framework.
@@ II-A Generation Tasks
Our evaluation framework assesses LLMs across three fundamental RTL code generation tasks, each representing a distinct level of complexity and scale. The smallest of tasks, in the sense of generation length, is Single-Line Completion (SLC), which focuses on the model’s ability to predict the next line of code given a partial context. This task closely resembles auto-completion scenarios, where engineers rely on coding assistants to streamline development.
@@
Finally, in Specification-to-RTL (S2R) the LLM must generate a complete implementation from a natural language hardware specification. This task does not provide a predefined module interface, so models must infer additional information from the specification, such as the module name, the correct ports and parameter types. This mimics the process of translating human-readable specifications into working hardware descriptions. Given their fundamental differences in complexity and nature, results for these three approaches are reported separately.
@@ II-B Design Goals
TuRTLe defines five design goals to evaluate LLM-generated RTL code, ensuring a comprehensive assessment of correctness, feasibility, and efficiency within the hardware design workflow. These goals are aligned with the supported tasks as shown in Table I.
Line-level Contextual Accuracy (LCA) measure the ability of the models to reproduce line completions in accordance to a given reference answer. This goal aligns with the SLC task, and partially evaluates whether the predicted line is both syntactically and functionally correct.
The four remaining goals can all be applied to both MC and S2R tasks. Syntax Correctness (STX) ensures that the generated HDL code complies with the language’s grammar rules and can be processed by standard HDL analyzers, synthesizers, and simulators [31]. The correctness of generated designs is verified by tools that check for syntax errors, missing constructs, and structural inconsistencies. Functional Correctness (FNC) ensures that the generated HDL code exhibits the expected behavior as specified in the design requirements (prompt) [32, 33]. FNC is verified through behavioral simulation, where the generated design is executed alongside a predefined testbench in an HDL simulator. While this approach provides a practical verification method, it relies on the assumption that testbenches accurately capture functional discrepancies.
Synthesizability (SYN) assesses whether the generated HDL code can be successfully synthesized into a gate-level netlist using a synthesis tool. In practical hardware design, these tools support only a subset of HDL constructs, making synthesizability a key criterion for real-world applicability. SYN is verified by processing the generated RTL code using a synthesis toolchain, ensuring that the design is implementable in hardware. This is critical for models that generate RTL intended for manufacturable chip designs.
Finally, Post-Synthesis Quality (PSQ) evaluates the implementation efficiency of the synthesized design based on Power, Performance, and Area (PPA). PPA metrics are widely used in chip design optimization to compare different implementations and ensure that a design meets key hardware constraints [34]. In this work, we use PPA to compare LLM-generated designs with human-crafted reference implementations. We set up a unified synthesis pipeline based on OpenLANE [35] to extract PPA from pairs of implementations of the same specification, one generated by an LLM and one by a human. Then, we extract area and power directly from the PPA report and evaluate performance based on maximum delay, defined as the difference between the clock period and the worst slack reported by static timing analysis.
TABLE I: Design goals and metrics for benchmarking RTL generation. The use of open-source tools ensures accessibility and reproducibility.
@@
SYN OpenLANE
PSQ PPA-Score OpenLANE
@@ II-C Measures and Metrics
To evaluate the five goals described above (LCA, STX, FNC, SYN and PSQ) we use different tools and measures tailored to each case. For LCA, prior work [28] has explored both exact and Fuzzy Matching (FM) techniques. Exact Matching (EM) determines whether the generated line precisely matches the reference HDL line in the dataset, while FM relies on edit distance and semantic similarity metrics. Our experimentation shows both metrics are strongly correlated, which is why we report only the more precise EM.
For MC and S2R tasks, we set up an evaluation pipeline that sequentially tests STX, FNC, SYN, and PSQ. STX is evaluated by compiling a design along with its testbench using Icarus Verilog [36] and checking for errors. If no errors are issued at compile time, FNC is evaluated by running the simulation executable generated by the compiler and checking if the testbench passes. Functionally correct codes are tested for SYN and PSQ, by synthesizing the design with OpenLANE [35]. For the LCA, STX, FNC, and SYN goals, we report the same metric Pass@$1$. This is a Pass@$k$ metric [37, 38], which measures the probability that at least one of $k$ generated solutions passes the corresponding test criteria. We set $k=1$ to focus on the realistic requirement of generating the correct result on the first try.
For designs that successfully pass SYN, we propose a novel evaluation metric called the PPA-score to assess post-synthesis design quality. This metric measures the performance of LLM-generated synthesizable code relative to a golden, human-crafted reference. The PPA-score is derived by extracting post-synthesis PPA metrics from the OpenLANE [35] reports. The designs are synthesized using the open-source SKY130A PDK [39] and constrained with a $10ns$ delay, which was empirically chosen considering the complexity of the designs within the tested benchmarks and the selected technology node [^1] Note that TuRTLe enables to change design constraints seamlessly through its configuration files.. Notice the framework and the analysis results are technology and constraints independent since identical settings are applied to both human- and LLM-generated designs, and the main goal remains to compare LLM performance against human baselines.
@@
Given a MC or S2R benchmark with $n$ problems, the TuRTLe framework generates $m$ candidate solutions for each. Then, each generation is processed sequentially through the evaluation pipeline: STX, FNC, SYN, PSQ. Each stage only processes results that passed the previous one, and failures in previous stages are reported as automatic fails in the next ones, creating a cascade score where $STX\geq FNC\geq SYN$. It could happen that $SYN<PSQ$ because we contemplate the possibility of having PSQ better than the human reference.
STX, FNC, and SYN are binary evaluations, pass or fail, which are aggregated using Pass@$k$. However, PSQ requires not only to aggregate real values that have to be analyzed in comparison with the reference PPA of the golden solution, but also to take into account that models that produce more synthetizable code can be evaluated on a larger set of problems, thus increasing the challenge and confidence in the results obtained. To achieve this, we let $p_{i,j}$ represent the PPA metric (power, performance, or area) from the LLM for candidate $j$ of problem $i$. The PPA-score is defined as the average of scores for each generation (represented as $\hat{p}_{i,j}$) computed according to the following steps:
@@
    Each $p_{i,j}$ is compared against the corresponding PPA value $g_{i}$ of the golden solution. For that, instead of aggregating $p_{i,j}$ we compute $p_{i,j}/g_{i}\in(0,+\infty)$.
@@
    For generations that do not pass STX, FNC and SYN evaluations, $p_{i,j}$ cannot be computed. As a result, we set a failure value of $2\cdot g_{i}$ (e.g., producing a design two times bigger than the human reference in the case of the area metric). This approach also clips to zero the score of designs that pass the previous evaluations but perform worse than this threshold.
@@
The score $\hat{p}_{i,j}$ of generation $j$ for problem $i$ is defined as
$\hat{p}_{i,j}=\begin{cases}2-\min(p_{i,j}/g_{i},2)&\text{if $p_{i,j}$ exists}\\ 0&\text{otherwise}\end{cases}$
which has the following interpretation:
  • $\hat{p}_{i,j}=0$ : means that the generation has not passed the previous stages, or that it requires twice or more the area, the power, or the delay (performance), when compared to the human reference.
  • $\hat{p}_{i,j}=1$ : are designs with an area, power, or performance equal to that of the human reference.
  • $\hat{p}_{i,j}=2$ : can only be obtained by chips which occupy no space, execute in no time, or consume no energy.
The final formula, considering all generations of an LLM for a given benchmark, is just the average of these scores:
$x\text{-score}=\left[\frac{1}{n\cdot m}\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{m}\hat{p}_{i,j}\right]\cdot 100\;\;[\%],$ (1)
@@
Since the framework includes results across multiple benchmarks, these need to be aggregated. A weighted average is used, considering the significant differences in individual benchmarks sizes (see Table II).